OpenStudy (anonymous):

A triangular prism has a volume of 64 yd3. If each dimension of the prism is reduced by half, what is the new volume of the prism?

OpenStudy (anonymous):

A triangular prism has a volume of\[V=\frac{1}{2}bhl.\]If we reduce each dimension by half, then\[V_{\text{new}}=\frac{1}{2}\frac{b}{2}\frac{h}{2}\frac{l}{2}=\frac{1}{16}bhl.\]What do you think?

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

I'm not sure. I think it could be something like 8 or 16...?

OpenStudy (anonymous):

The answer is that the volume is reduced by a factor of \(8\).

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

What is the answer ,ANYBODY !?